OpenStudy (k.binks):

What is the equation of the ellipse with vertices (5, 0), (-5, 0) and co-vertices (0, 4), (0, -4)?

OpenStudy (danjs):

standard form when centered at the origin \[\frac{ x^2 }{ a^2 }+ \frac{ y^2 }{ b^2 } = 1\]

OpenStudy (danjs):

intersects x axis at plus or minus a y axis at plus or minus b

OpenStudy (k.binks):

I'm sorry, I still don't understand. This is the first time I've done an assignment on ellipses. Could you show me step by step how to do this?

OpenStudy (loser66):

 plug into the standard form above. done

OpenStudy (loser66):

still not get?? or doubt because how easy it is??? hahaha.... it is really easy!!

OpenStudy (k.binks):

I am really, really bad at math. I can never tell if I'm doing it right when I plug in equations. x^2/25 +y^2/16= 1?

OpenStudy (loser66):

yes, that is it!!

OpenStudy (k.binks):

Great, thankyou!